OpenStudy (kearsten):

medal and fan solve the system using elimination -2x+3y=9 and 2x-2y=-4? show steps if possible

OpenStudy (anonymous):

You use elimination when there are or you are able to get rid of a variable by doing an operation between the two equations. In this case, one equation has a -2x, and the other has a 2x. What operation do you do to get rid of the xs?

OpenStudy (kearsten):

subtraction?

OpenStudy (anonymous):

Not quite. Lets call 2x-2y = -4 equation 1, and -2x + 3y = 9 equation 2. If you mean that we subtract equation 1 from equation 2, you'll get 4x-5y=-13. There are too many variables to solve the system by subtraction. What other operation can we do?

OpenStudy (kearsten):

addition?

OpenStudy (anonymous):

Yes. Now, what do we get when you add the two equations?

OpenStudy (kearsten):

9y=5?

OpenStudy (anonymous):

Almost. There's no 9 infront of the y.-2y +3y =y

OpenStudy (anonymous):

So, we know the value of one variable. What can we do with that variable that will give us the other variable? Consider that we still have those two equations from before.

OpenStudy (kearsten):

insert it to equation 1?

OpenStudy (anonymous):

Do you still need help with the question? I want to help out as well

OpenStudy (anonymous):

Right. What is the value of x when you put it in equation 1?

OpenStudy (kearsten):

2x-2=-4,

OpenStudy (kearsten):

x=1?

OpenStudy (anonymous):

Remember that we found that y = 5. Lets use equation 2. It doesn't matter which equation we use, so pick eithier that you'd like. equation 2 was: 2x-2y = -4. We know that y = 5 because we just found that. 2x - 2(5) = -4 2x-10 = -4 Now we need to isolate x. How can we do that?

OpenStudy (kearsten):

divide -4 by 2 and +10?

OpenStudy (wolf1728):

If you are willing to start again here's the equations: A) -2x+3y=9 B) 2x-2y=-4 We add the 2 equations to get y = 5 Now to get the value of x, put the value of y into equation A or equation B

OpenStudy (anonymous):

If you want to divide first , you need to divide the 10 by 2 too, to get 5. You can move the five as well. Doing this, you'll find that x = 3.

OpenStudy (anonymous):

If you want to check your answer, put x = 3 and y = 5 into equation 1. If left side equals right side, you have the right answers.

OpenStudy (kearsten):

thank you Hari_Seldon

OpenStudy (wolf1728):

kearsten - want me to go over that with you?

OpenStudy (kearsten):

no, thanks. I think I got it but thanks for your help wolf1728

OpenStudy (wolf1728):

Okay - good luck :-)

OpenStudy (kearsten):

i suck at math and you guys really helped

OpenStudy (wolf1728):

glad we could help